The velocity components in a two dimensional flow are :
U = 8x^2 y - 8/3 y^3 And
V = -8xy^3 + 8/3 x^3.
Show that these velocity components represents a possible case of an irrotational flow.
1
Expert's answer
2020-12-16T10:27:22-0500
"u=8x^2y-\dfrac{8}{3}y^3, v=-8xy^3+\dfrac{8}{3}x^3"


The fluid is irrotational, if "\dfrac{\partial u}{\partial y}=\dfrac{\partial v}{\partial x}."



"\dfrac{\partial u}{\partial y}=8x^2-8y^2, \dfrac{\partial v}{\partial x}=-8y^3+8x^2"

"\dfrac{\partial u}{\partial y}\not=\dfrac{\partial v}{\partial x}"


Therefore, it is not a possible irrotational flow.


If

"u=8x^2y-\dfrac{8}{3}y^3, v=-8xy^2+\dfrac{8}{3}x^3"

The fluid is irrotational, if "\dfrac{\partial u}{\partial y}=\dfrac{\partial v}{\partial x}."



"\dfrac{\partial u}{\partial y}=8x^2-8y^2, \dfrac{\partial v}{\partial x}=-8y^2+8x^2"

"\dfrac{\partial u}{\partial y}=\dfrac{\partial v}{\partial x}"


Therefore, these velocity components represents a possible case of an irrotational flow.